OpenStudy (lgbasallote):

when hydrogen atom jumps from the n=4 state to the n=2 state, a photon is emitted. Calculate \(\Delta') E for the transition. can someone show me the solution for this?

OpenStudy (anonymous):

First off, hydrogen atoms have s-orbitals and one lone electron. This makes it a quite a bit easier, when talking quantum states. Now you're talking about a drop of two states, unfortunately the states you are talking in are not as simple as you might think, look: http://upload.wikimedia.org/wikipedia/commons/thumb/e/e7/Hydrogen_Density_Plots.png/1126px-Hydrogen_Density_Plots.png This can be described by a set of 4 numbers, called the quantum state: \[\large( n , \ell , m_\ell , m_s)\] As anybody who has dabbled in quantum physics can tell you though: the more you know about the position, the less you can tell about it's properties -- the more you know about the properties, the less you know about the position. Thankfully here, you can drop the properties and all the complexities from the problem itself because the quantum system you have here is very simple. Phew! So... \[Energy \approx \large -13.6eV(\frac{1}{n_f^2}-\frac{1}{n_i^2})=-13.6eV\frac{1}{(4)^2}-\frac{1}{(2)^2}\] 1 eV = 1.6022 × 10^-19 Joules Can you figure it out from here? Post back if you do I'd like to know if it's the correct answer. :-)

OpenStudy (anonymous):

Hmm, I'm trying to see if I have initial and final switched. In any case, here ya go: http://cas.sdss.org/dr6/en/proj/advanced/spectraltypes/energylevels.asp

OpenStudy (lgbasallote):

wait...dont you multiply by -2.18 x10^-18?

OpenStudy (anonymous):

-2.18 x10\(^{-18}\) = -13.6 eV *\(\huge\frac{1.6022 \cdot10^{-19} J}{1 eV}\)

OpenStudy (anonymous):

Units cancel out ya know :-P

OpenStudy (anonymous):

You're left with an answer in Joules.

OpenStudy (lgbasallote):

but i came up with -5.45 x 10^-19 j using -2.18 x 10^-18

OpenStudy (anonymous):

Well conversion factor rather, but whatever

OpenStudy (anonymous):

Check you parenthesis if you're using a calculator, TI-84's for example don't handle well with arithmetic and scientific notation. The -2.18 number is validated by the equation show just above. I had to type it in myself when I was like, "what the heck is that number?!"

OpenStudy (anonymous):

-5.45 x 10\(^{-19}\) J doesn't sound right at all... Not sure where that number came from

OpenStudy (anonymous):

What that your final answer? o_O

OpenStudy (lgbasallote):

let me post my solution..wait a minute..

OpenStudy (lgbasallote):

oh lol i see...i forgot to square *facepalm*

OpenStudy (lgbasallote):

i'll try re-solving

OpenStudy (anonymous):

A release of about 4 x 10\(^{-19}\) Joules is what I get, can you confirm?

OpenStudy (lgbasallote):

yup that's what i get too

OpenStudy (lgbasallote):

so that's \(\Delta\) E?

OpenStudy (anonymous):

That's the change in energy yes, make sure you not it's energy being released by the system (the atom) out into the environment (via a photon)

OpenStudy (anonymous):

make sure you note* sry been a long day